OpenStudy (anonymous):

substitution for 4x-6y =6 and x + 7y = -7

OpenStudy (anonymous):

4x-6y=6 x=(6+6y)/4 ...i Also, x+7y=-7 (6+6y)/4 +7y=-7 [from i] Now, solve for y

OpenStudy (anonymous):

Can u?